A train travels x miles in 2 hours.
Write down an expression , in terms of x , for the average train speed

Answer by Theo(13342)

rate * time = distance
solve for rate to get:
rate = distance / time
let y = rate and x = distance and time = 2 and the formula becomes:
y = x / 2
that's your formula for average train speed in terms of x.
that equation can be graphed as shown below.
x is the distance.
when x = 100, y = 100/2 = 50 miles per hour
when x = 150, y = 150/2 = 75 miles per hour.
time in this equation is always 2 hours.